A prismatic bar with rectangular cross-section (25mmX45mm) and length 3m is subjected to an axial tensile force of 70KN. The measured deformation of the bar is 1.2mm. Calculate the tensile stress and strain in the bar.


1
Expert's answer
2020-09-06T17:17:27-0400

The stress

"\sigma=\frac{F}{A}=\rm\frac{70\;000\: N}{0.025\: m\times 0.045\: m}=6.2\times 10^7\: Pa"

The strain

"\varepsilon=\frac{\Delta L}{L_0}=\rm\frac{0.0012\: mm}{3 \: m}=0.0004"